OpenStudy (anonymous):

What are the solutions to x^2+4x+4=9? help?

OpenStudy (anonymous):

(x+2)^2 =9 x+2 = +-3

OpenStudy (anonymous):

x=1 and x= -5

OpenStudy (anonymous):

quadratic equation

OpenStudy (anonymous):

(x+5)(x-1)=0 x=1,-5

OpenStudy (anonymous):

its a quadratic equation ,(x+5)(x-1)=0 and so u get 1,-5.